### v2

@@ 3.2 Fragility, robustness and default cascades
With the given assumption, we have $D_{t}^{(0)}=D_{t-}$. Then it can be inductively shown that $D_{t}^{(N)}=D_{t-}$ for any $N\geq 1$. Therefore, $D_{t}=D_{t-}$. ∎
Next, we are going to prove another qualitative property of the jump sizes for the minimal solutions. We start with the following lemma, which says that, with probability $1$, there can be at most one default cascade triggered at a specific time.
Lemma 3.10.
Under the same assumptions as in Lemma 3.5, we have
$\mathbb{P}\Bigg{[}\Big{|}\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t-}\}}=0\Big{\}}\Big{|}\leq 1,\,\forall t\geq 0\Bigg{]}=1.$
Proof.
With probability one, a non-degenerate two-dimensional Brownian motion will not hit a given point that is not its starting point. It now suffices to note that
$\Bigg{\{}\exists t\geq 0\mathrel{\mathop{\ordinarycolon}}\,\Big{|}\Big{\{}v\in V\mathrel{\mathop{\ordinarycolon}}x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t-}\}}=0\Big{\}}\Big{|}\geq 2\Bigg{\}}$
$\subset$ $\bigcup_{u\in V}\bigcup_{v\in V}\bigcup_{N_{1}\subset N_{G}^{-}(u)}\bigcup_{N_{2}\subset N_{G}^{-}(v)}\Big{\{}\exists t\geq 0\mathrel{\mathop{\ordinarycolon}}\,x_{u}+B_{u}(t)=\sum_{w\in N_{1}}c_{wu},\quad\,x_{v}+B_{v}(t)=\sum_{w\in N_{2}}c_{wv}\Big{\}},$
the latter of which is a countable union of events of probability 0. ∎
Proposition 3.11.
Let the configuration $(G,c,x,B)$ drive robust systems and let $(X,D)$ be the minimal solution. Then
$\mathbb{P}[|D_{t}\setminus D_{t-}|<\infty,\quad\forall t\geq 0]=1.$
Proof.
Consider any $t\geq 0$ that satisfies $D_{t-}\subsetneq D_{t}$. By the previous lemma, there exists a unique $v_{0}\notin D_{t-}$ such that
$x_{v_{0}}+B_{v_{0}}(t)-\sum_{u\in N_{G}^{-}(v_{0})}c_{uv_{0}}\mathbf{1}_{\{u\in D_{t-}\}}=0.$
Recall from the proof of Proposition 3.7 that
$D_{t}\setminus D_{t-}\subset F_{t}\mathrel{\mathop{\ordinarycolon}}=\left\{v\in V\mathrel{\mathop{\ordinarycolon}}0\leq x_{v}+B_{v}(t)\leq w_{v}\right\},$
where each weakly connected component of $F_{t}$ is finite almost surely due to robustness.Let $F_{t}=\bigcup_{n}F_{t}^{n}$ be the decomposition into weakly connected components, and let $n_{0}\in\mathbb{N}$ be such that $v_{0}\in F_{t}^{n_{0}}$.
We claim that $D_{t}\subset D_{t-}\cup F_{t}^{n_{0}}$. First, by the definitions we immediately have $D_{t}^{(0)}=D_{t-}\cup\{v_{0}\}\subset D_{t-}\cup F_{t}^{n_{0}}$. Suppose that we already have $D_{t}^{(N)}\subset D_{t-}\cup F_{t}^{n_{0}}$. Then for any $v\in V$ such that
$x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(n)}\}}\leq 0,$
either $v\in D_{t}^{(N)}\subset D_{t-}\cup F_{t}^{n_{0}}$, or
$x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}^{(N-1)}\}}>0$
but there exists $u\in N_{G}^{-}(v)$ such that $u\in D_{t}^{(N)}\setminus D_{t}^{(N-1)}\subset F_{t}^{n_{0}}$. For the latter case, since the above two inequalities imply $v\in F_{t}$, the existence of such $u$ entails $v\in F_{t}^{n_{0}}$ by the maximal connectedness of $F_{t}^{n_{0}}$. This proves $D_{t}^{(N+1)}\subset D_{t-}\cup F_{t}^{n_{0}}$. Hence, it can be shown inductively that $D_{t}^{(N)}\subset D_{t-}\cup F_{t}^{n_{0}}$ for any $N\geq 0$ and therefore $D_{t}\subset D_{t-}\cup F_{t}^{n_{0}}$, or equivalently, $D_{t}\setminus D_{t-}\subset F_{t}^{n_{0}}$. As $F_{t}^{n_{0}}$ is finite, we also find that $D_{t}\setminus D_{t-}$ must be finite almost surely. ∎
Remark 3.12.
The theorem provides another theoretical support for the findings of [2]: compared to fully connected or densely connected networks, in some circumstances, sparsely connected networks are more resilient to systemic risk. Our analysis reveals the mechanism behind this phenomenon—in locally connected networks, the robustness condition ensures that contagion remains confined to finite connected components of fragile institutions. Well-capitalized banks outside these vulnerable clusters act as natural “firebreaks”, preventing cascade propagation to the broader part of the network. This containment property is formalized through our robustness condition, which guarantees that infinite percolation of fragile institutions cannot occur. In contrast, such safety barriers are clearly absent in fully connected networks, and default cascades can potentially bring down a significant proportion of institutions throughout the whole network, as demonstrated by the mean-field model proposed in [31].
@@
Let the configuration $(G,c,x,B)$ drive $\delta$-robust systems, for some $\delta>0$. Let $(X,D)$ be a physical solution. Then it is the minimal solution.
@@
Let $(\underline{X},\underline{D})$ be the minimal solution, and let $t_{0}\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0,D_{t}\neq\underline{D}_{t}\}$. Suppose that, for contradiction, $t_{0}<\infty$. Then $D_{s}=\underline{D}_{s}$ and also $X_{v}(s)=\underline{X}_{v}(s)$ for any $s\in[0,t_{0})$ and any $v\in V$. By continuity of $X_{v}$ this implies that $D_{t_{0}}^{(0)}=\underline{D}_{t_{0}}^{(0)}$. Iterating the physical condition then gives $D_{t_{0}}=\underline{D}_{t_{0}}$. Now, it follows from the definition of $t_{0}$ that, for any $\delta>0$, there must exist $v_{\delta}^{(0)}\in D_{t_{0}+\delta}\setminus\underline{D}_{t_{0}+\delta}$. Then there must exist $v_{\delta}^{(1)}\in N_{G}^{-}(v_{\delta}^{(0)})\cap(D_{t_{0}+\delta}\setminus\underline{D}_{t_{0}+\delta}$) that defaults before $v_{\delta}^{(0)}$ in the system $(X,D)$. Iterating this argument, we get a sequence $(v_{\delta}^{(n)})_{n\geq 0}\subset D_{t_{0}+\delta}\setminus\underline{D}_{t_{0}+\delta}$ such that $v_{\delta}^{(n+1)}\in N_{G}^{-}(v_{\delta}^{(n)})$ and that $v_{\delta}^{(n+1)}$ defaults before $v_{\delta}^{(n)}$ in the system $(X,D)$. The existence of such a sequence implies that $\bigcup_{t\in[t_{0},t_{0}+\delta]}F_{t}$ has an infinitely large weakly connected component, which is a contradiction to the $\delta$-robustness assumption. As a result, we must have $t_{0}=\infty$ and thus $D_{t}=\underline{D}_{t}$ for any $t\geq 0$. That is, $(X,D)$ is the minimal solution. ∎
@@ 3.3 Default trees and recovery of locality
In this subsection, we provide a more quantitative description of the interaction among the particles that are distant away from each other. This is analyzed via the growth estimate of default trees, defined in the following.
@@
Let the configuration $(G,c,x,B)$ drive $\delta$-robust systems and let $(X,D)$ be the associated physical solution. For any vertex $v_{0}\in V$ such that $\tau_{v_{0}}<\infty$, the default tree $\mathcal{T}(G,v_{0})$ rooted at $v_{0}$ is a subgraph of $G$ defined recursively as follows:
@@
Let the configuration $(G,c,x,B)$ drive $\delta$-robust systems for some $\delta>0$, and let a sequence of configurations $(G_{n},c^{n},x^{n},B^{n})_{n\geq 1}$ converges to $(G,c,x,B)$ locally in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$ almost surely as $n\to\infty$. Then with probability one, for any fixed $v_{0}\in V$ satisfying either:
@@
$0\leq x_{v}^{n}+B_{v}(\tau_{v_{0}}^{n})\leq\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}.$
Taking $n\to\infty$ gives
$0\leq x_{v}+B_{v}(\tau)\leq\sum_{u\in N_{G}^{-}(v)}c_{uv}$
@@
$\mathcal{L}_{m}\mathrel{\mathop{\ordinarycolon}}=\{l\in\mathbb{N}\mathrel{\mathop{\ordinarycolon}}\,\exists m^{\prime}\in\{m+1,...,M\}\,\exists l^{\prime}\in\mathcal{L}_{m^{\prime}},F_{[m\delta,(m+1)\delta]}^{l}\cap B_{G}(F_{[m^{\prime}\delta,(m^{\prime}+1)\delta]}^{l^{\prime}},1)\neq\emptyset\}.$
@@
Step 3. As an intermediate step, we prove that $\tau_{v_{0}}\geq\limsup_{n\to\infty}\tau_{v_{0}}^{n}$. Since the desired inequality is trivial if $\tau_{v_{0}}=\infty$, we focus on the case in which $\tau_{v_{0}}<\infty$. As $\mathcal{T}(G,v_{0})$ is a finite tree, we can use a bottom-up induction to show that $\tau_{v}\geq\limsup_{n\to\infty}\tau_{v}^{n}$ for all $v\in\mathcal{T}(G,v_{0})$. For the base case, suppose $v$ is a leaf node of $\mathcal{T}(G,v_{0})$. Then necessarily $X_{v}(\tau_{v})=x_{v}+B_{v}(\tau_{v})=0$. With probability $1$, for any $\Delta>0$, there exists $t_{\Delta}\in(\tau_{v},\tau_{v}+\Delta)$ such that $x_{v}+B_{v}(t_{\Delta})<0$. For all sufficiently large $n$, it must hold that $x_{v}^{n}+B_{v}^{n}(t_{\Delta})<0$ and thus $\limsup_{n\to\infty}\tau_{v}^{n}\leq\tau_{v}+\Delta$. Letting $\Delta\downarrow 0$ gives $\tau_{v}\geq\limsup_{n\to\infty}\tau_{v}^{n}$. For the inductive step, we assume that $v$ is a node such that any of its children $u$ in $\mathcal{T}(G,v_{0})$ satisfies $\tau_{u}\geq\limsup_{n\to\infty}\tau_{u}^{n}$. Similarly, as
$X_{v}(\tau_{v})=x_{v}+B_{v}(\tau_{v})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}\leq 0,$
with probability $1$, for any $\Delta>0$, there exists $t_{\Delta}\in(\tau_{v},\tau_{v}+\Delta)$ such that
$x_{v}+B_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}<0.$
@@
$x_{v}^{n}+B_{v}^{n}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in D_{t_{\Delta}}^{n}\}}<0,$
@@
$\sup_{s\in[m\delta,(m+1)\delta]}(x_{v}+B_{v}(s))<0\quad\text{or}\quad\inf_{s\in[m\delta,(m+1)\delta]}(x_{v}+B_{v}(s))>\sum_{u\in N_{G}^{-}(v)}c_{uv}.$
Now we can take $N_{1}\geq N_{0}$ such that for all $n\geq N_{1}$:
$\sup_{s\in[m\delta,(m+1)\delta]}(x_{v}^{n}+B_{v}^{n}(s))<0\quad\text{or}\quad\inf_{s\in[m\delta,(m+1)\delta]}(x_{v}^{n}+B_{v}^{n}(s))>\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}$
for all $v\in B_{G}(V_{0},1)$ such that $v\notin F_{[m\delta,(m+1)\delta]}$, and all $m\in\{0,1,...,M\}$. In particular, for any $t\in[0,(M+1)\delta]$, any $v\in B_{G}(V_{0},1)$ and any $n\geq N_{1}$, $v$ being fragile subject to $(G_{n},c^{n},x^{n},B^{n})$ at time $t\in[m\delta,(m+1)\delta]$ implies that $v\in F_{[m\delta,(m+1)\delta]}$.
@@ 3.4 Convergence of physical solutions
Below, we present and prove a result that is more general than Theorem 1.9. Here, we treat the driving noises $(B_{v})_{v\in V}$, which are not necessarily standard Brownian motions, as part of the inputs to the equations (1.1) and (1.2). It turns out that the only technical assumptions that are essential are the $\delta$-robustness of the limiting configuration $(G,c,x,B)$ and the downward crossing property on the driving noises defined below:
Definition 3.18.
A stochastic process $b_{t}$ is said to satisfy the downward crossing property if, for any fixed stopping time $\tau$,
$\mathbb{P}\Big{[}\forall h>0,\,\inf_{s\in[0,h]}(b_{\tau+s}-b_{\tau})<0\Big{]}=1.$
The functional that takes a càdlàg process to its zero-hitting time is continuous at those processes satisfying the downward crossing property (see e.g. [16, Proposition 5.8]). Formally speaking, any càdlàg stochastic process with a Brownian component satisfies the above downward crossing property.
Theorem 3.19.
Assume that $(G_{n},c^{n},x^{n},B^{n})$ drives robust systems for each $n\geq 1$, $(G,c,x,B)$ drives $\delta$-robust systems for some $\delta>0$, and $B_{v}$ satisfies the downward crossing property for any $v\in G$. If $\mathcal{L}(G_{n},c^{n},x^{n},B^{n})\to\mathcal{L}(G,c,x,B)$ in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}])$ as $n\to\infty$, then $\mathcal{L}(G_{n},\underline{X}^{n},\underline{D}^{n})\to\mathcal{L}(G,\underline{X},\underline{D})$ in $\mathcal{P}(\mathcal{G}_{*}[\mathcal{D}^{2}])$.
Proof.
Step 1. We start by showing that the sequence $\{\mathcal{L}(G_{n},c^{n},x^{n},B^{n},\underline{X}^{n},\underline{D}^{n})\}_{n\geq 1}$ is tight in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}\times\mathcal{D}^{2}])$. By the Skorokhod Representation Theorem, we can assume without loss of generality that $(G_{n},c^{n},x^{n},B^{n})$ converges almost surely to $(G,c,x,B)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$ as $n\to\infty$.
Fix any $\varepsilon\in(0,1)$. Since local convergence holds almost surely, we can find a compact subset $\mathcal{K}_{0}^{\varepsilon}$ of $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}]$ together with a sequence $(M_{m})_{m\geq 1}$ of positive real numbers such that
$\inf_{n}\mathbb{P}\Big{[}(G_{n},c^{n},x^{n},B^{n})\in\mathcal{K}_{0}^{\varepsilon},\,\max_{v\in B_{G}(o,m)}\sum_{u\in N_{G}^{-}(v)}c_{uv}\leq M_{m}\Big{]}\geq 1-\frac{\varepsilon}{2}.$
@@
$\inf_{n}\mathbb{P}\Big{[}(G_{n},c^{n},x^{n},B^{n})\in\mathcal{K}_{0}^{\varepsilon},\,\underline{L}_{v}^{n}\in\mathcal{M}(M_{m})\text{ and }\underline{D}_{v}^{n}\in\mathcal{M}(1),\,\forall v\in B_{G_{n}}(o_{n},m)\,\forall m\geq 1\Big{]}\geq 1-\varepsilon,$
which implies the tightness of $\{\mathcal{L}(G_{n},c^{n},x^{n},B^{n},\underline{L}^{n},\underline{D}^{n})\}_{n\geq 1}$ in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}\times\mathcal{D}^{2}])$ by Theorem 2.8 and Lemma 2.4. As
$\underline{X}_{v}^{n}=x_{v}^{n}+B_{v}^{n}-\underline{L}_{v}^{n},\quad v\in V_{n},$
we obtain that $\{\mathcal{L}(G_{n},c^{n},x^{n},B^{n},\underline{X}^{n},\underline{D}^{n})\}_{n\geq 1}$ is tight in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}\times\mathcal{D}^{2}])$ by combining Lemma 2.3 Lemma A.1 and Lemma A.5.
Step 2. By the Skorokhod Representation Theorem, we can find a $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}\times\mathcal{D}^{2}]$-valued random element $(G,c,x,B,X,D)$ such that $(G_{n},c^{n},x^{n},B^{n},\underline{X}^{n},\underline{D}^{n})\stackrel{{\scriptstyle a.s.}}{{\to}}(G,c,x,B,X,D)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{C}\times\mathcal{D}^{2}]$ as $n\to\infty$. In this step, we verify that $(X,D)$ is a solution to equations (1.1) and (1.2). As
$\underline{X}_{v}^{n}(t)$ $=x_{v}^{n}+B_{v}^{n}(t)-\sum_{u\in N_{G_{n}}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in D_{t}^{n}\}},\quad v\in V_{n},$
by Theorem A.2, for all $t\in[-1,\infty)$ such that $t$ is a continuity point of $(X_{v}(\cdot),\mathbf{1}_{\{v\in D_{\cdot}\}})$ for all $v\in V$ (the set of such $t$ is co-countable and thus dense in $1,\infty)$), it holds that
$X_{v}(t)$ $=x_{v}+B_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}},\quad v\in V.$
@@
by Corollary A.3. To prove $\tau_{v}=\lim_{n\to\infty}\underline{\tau}_{v}^{n}$, we can use the same argument as in the proof of [16, Proposition 5.8], as the limit process $X_{v}$ satisfies the downward crossing property.

### v3

@@ 3.2 Fragility, robustness and default cascades
With the given assumption, we have $D_{t}^{(0)}=D_{t-}$. Then it can be inductively shown that $D_{t}^{(N)}=D_{t-}$ for any $N\geq 1$. Therefore, $D_{t}=D_{t-}$ by the physicality condition. ∎
@@
Let the configuration $(G,c,x,Z)$ drive $\delta$-robust systems, for some $\delta>0$. Let $(X,D)$ be a physical solution. Then it is the minimal solution.
@@
Let $(\underline{X},\underline{D})$ be the minimal solution, and let $t_{0}\mathrel{\mathop{\ordinarycolon}}=\inf\{t\geq 0,D_{t}\neq\underline{D}_{t}\}$. Suppose that, for contradiction, $t_{0}<\infty$. Then $D_{s}=\underline{D}_{s}$ and also $X_{v}(s)=\underline{X}_{v}(s)$ for any $s\in[0,t_{0})$ and any $v\in V$. By continuity of $X_{v}$ this implies that $D_{t_{0}}^{(0)}=\underline{D}_{t_{0}}^{(0)}$. Iterating the physical condition then gives $D_{t_{0}}=\underline{D}_{t_{0}}$. Now, it follows from the definition of $t_{0}$ that, for any $\delta>0$, there must exist $t_{\delta}\in(t_{0},t_{0}+\delta)$ and $v_{\delta}^{(0)}\in D_{t_{\delta}}\setminus\underline{D}_{t_{\delta}}$. Let $(\tau,k)\mathrel{\mathop{\ordinarycolon}}=(\tau_{v_{\delta}^{(0)}},k_{v_{\delta}^{(0)}})$, then we must have $v_{\delta}^{(0)}\in D_{\tau}^{(k)}\setminus\underline{D}_{\tau}$, which implies
$x_{v_{\delta}^{(0)}}+Z_{v_{\delta}^{(0)}}(\tau)-\sum_{u\in N_{G}^{-}({v_{\delta}^{(0)}})}c_{u{v_{\delta}^{(0)}}}\mathbf{1}_{\{u\in D_{\tau}^{(k-1)}\}}\leq 0$ (3.3)
while $x_{v_{\delta}^{(0)}}+Z_{v_{\delta}^{(0)}}(\tau)-\sum_{u\in N_{G}^{-}({v_{\delta}^{(0)}})}c_{u{v_{\delta}^{(0)}}}\mathbf{1}_{\{u\in\underline{D}_{\tau}\}}>0.$
Hence there must exist $v_{\delta}^{(1)}\in N_{G}^{-}(v_{\delta}^{(0)})\cap(D_{\tau}^{(k-1)}\setminus\underline{D}_{\tau}$). In other words, $v_{\delta}^{(1)}$ is an element of $N_{G}^{-}(v_{\delta}^{(0)})$ that defaults before $v_{\delta}^{(0)}$ in the system $(X,D)$. Moreover, Equations (LABEL:eq:UniqueMinEq1) imply that $v_{\delta}^{(0)}\in F_{\tau}\subset\bigcup_{t\in[t_{0},t_{0}+\delta]}F_{t}$. Iterating this argument, we get a sequence $(v_{\delta}^{(n)})_{n\geq 0}\subset D_{t_{\delta}}\setminus\underline{D}_{t_{\delta}}$ such that $v_{\delta}^{(n+1)}\in N_{G}^{-}(v_{\delta}^{(n)})$ and that $v_{\delta}^{(n+1)}$ defaults before $v_{\delta}^{(n)}$ in the system $(X,D)$. The existence of such a sequence implies that $\bigcup_{t\in[t_{0},t_{0}+\delta]}F_{t}$ has an infinitely large weakly connected component, which is a contradiction to the $\delta$-robustness assumption. As a result, we must have $t_{0}=\infty$ and thus $D_{t}=\underline{D}_{t}$ for any $t\geq 0$. That is, $(X,D)$ is the minimal solution. ∎
@@ 3.3 Default trees and recovery of locality
Before discussing the proof of Theorem 1.11, we need a more accurate description of the interaction among the particles that are potentially distant away from each other, which is of interest on its own. This is analyzed via the growth of default trees, defined in the following.
@@
Let $(X,D)$ be a physical solution associated with a problem configuration $(G,c,x,Z)$. For any vertex $v_{0}\in V$ such that $\tau_{v_{0}}<\infty$, the default tree $\mathcal{T}(G,v_{0})$ rooted at $v_{0}$ is a subgraph of $G$ defined recursively as follows:
@@
Let the configuration $(G,c,x,Z)$ drive $\delta$-robust systems for some $\delta>0$, and let a sequence of configurations $(G_{n},c^{n},x^{n},Z^{n})_{n\geq 1}$ converges to $(G,c,x,Z)$ locally in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$ almost surely as $n\to\infty$. Then with probability one, for any fixed $v_{0}\in V$ satisfying either:
@@
$x_{v_{0}}^{n}+Z_{v_{0}}^{n}(\tau_{v_{0}}^{n}-)\geq 0\quad\text{and}\quad x_{v_{0}}^{n}+Z_{v_{0}}^{n}(\tau_{v_{0}}^{n})\leq\sum_{u\in N_{G}^{-}(v_{0})}c_{uv_{0}}^{n}.$
Taking $n\to\infty$, applying Lemma B.5 and making use of the assumption that $Z_{v_{0}}$ does not jump upward (which implies $Z_{v_{0}}(\tau)\leq Z_{v_{0}}(\tau-)$), we obtain
$x_{v_{0}}+Z_{v_{0}}(\tau-)\geq 0\quad\text{and}\quad x_{v_{0}}+Z_{v_{0}}(\tau)\leq\sum_{u\in N_{G}^{-}(v_{0})}c_{uv_{0}}.$
@@
$\mathcal{L}_{m}\mathrel{\mathop{\ordinarycolon}}=\{l\in\mathbb{N}\mathrel{\mathop{\ordinarycolon}}\,\exists m^{\prime}\in\{m+1,.,M\}\,\exists l^{\prime}\in\mathcal{L}_{m^{\prime}},F_{[m\delta,(m+1)\delta]}^{l}\cap B_{G}(F_{[m^{\prime}\delta,(m^{\prime}+1)\delta]}^{l^{\prime}},1)\neq\emptyset\}.$
@@
Step 3. As an intermediate step, we prove that $\tau_{v_{0}}\geq\limsup_{n\to\infty}\tau_{v_{0}}^{n}$. Since the desired inequality is trivial if $\tau_{v_{0}}=\infty$, we focus on the case in which $\tau_{v_{0}}<\infty$. As $\mathcal{T}(G,v_{0})$ is a finite tree, we can use a bottom-up induction to show that $\tau_{v}\geq\limsup_{n\to\infty}\tau_{v}^{n}$ for all $v\in\mathcal{T}(G,v_{0})$. For the base case, suppose $v$ is a leaf node of $\mathcal{T}(G,v_{0})$. Then necessarily $X_{v}(\tau_{v})=x_{v}+Z_{v}(\tau_{v})\leq 0$. With probability $1$, for any $\Delta>0$, there exists $t_{\Delta}\in(\tau_{v},\tau_{v}+\Delta)$ ,which can be taken to be a continuity point of $Z_{v}(\cdot)$, such that $x_{v}+Z_{v}(t_{\Delta})<0$. For all sufficiently large $n$, it must hold that $x_{v}^{n}+Z_{v}^{n}(t_{\Delta})<0$ and thus $\limsup_{n\to\infty}\tau_{v}^{n}\leq\tau_{v}+\Delta$. Letting $\Delta\downarrow 0$ gives $\tau_{v}\geq\limsup_{n\to\infty}\tau_{v}^{n}$. For the inductive step, we assume that $v$ is a node such that any of its children $u$ in $\mathcal{T}(G,v_{0})$ satisfies $\tau_{u}\geq\limsup_{n\to\infty}\tau_{u}^{n}$. Similarly, as
$X_{v}(\tau_{v})=x_{v}+Z_{v}(\tau_{v})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}\leq 0,$
with probability $1$, for any $\Delta>0$, there exists $t_{\Delta}\in(\tau_{v},\tau_{v}+\Delta)$, that can be taken to be a continuity point of $Z_{v}(\cdot)$, such that
$x_{v}+Z_{v}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{\tau_{v}}\}}<0.$
@@
$x_{v}^{n}+Z_{v}^{n}(t_{\Delta})-\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in D_{t_{\Delta}}^{n}\}}<0,$
@@
$\sup_{s\in[m\delta,(m+1)\delta]}(x_{v}+Z_{v}(s))<0\quad\text{or}\quad\inf_{s\in[m\delta,(m+1)\delta]}(x_{v}+Z_{v}(s))>\sum_{u\in N_{G}^{-}(v)}c_{uv}.$
As $Z_{v}(\cdot)$ is càdlàg, there exists a $\delta_{v}\in(0,\delta)$ such that
$\sup_{s\in[m\delta-\delta_{v},(m+1)\delta+\delta_{v}]}(x_{v}+Z_{v}(s))<0\quad\text{or}\quad\inf_{s\in[m\delta-\delta_{v},(m+1)\delta+\delta_{v}]}(x_{v}+Z_{v}(s))>\sum_{u\in N_{G}^{-}(v)}c_{uv}.$
By Lemma B.6, we can take $N_{1}\geq N_{0}$ such that for all $n\geq N_{1}$:
$\sup_{s\in[m\delta,(m+1)\delta]}(x_{v}^{n}+Z_{v}^{n}(s))<0\quad\text{or}\quad\inf_{s\in[m\delta,(m+1)\delta]}(x_{v}^{n}+Z_{v}^{n}(s))>\sum_{u\in N_{G}^{-}(v)}c_{uv}^{n}$
for all $v\in B_{G}(V_{0},1)$ such that $v\notin F_{[m\delta,(m+1)\delta]}$, and all $m\in\{0,1,...,M\}$. In other words, for any $t\in[0,(M+1)\delta]$, any $v\in B_{G}(V_{0},1)$ and any $n\geq N_{1}$, $v$ being fragile subject to $(G_{n},c^{n},x^{n},Z^{n})$ at time $t\in[m\delta,(m+1)\delta]$ implies that $v\in F_{[m\delta,(m+1)\delta]}$.
@@ 3.4 Convergence of physical solutions
Below, we present and prove a result that is more general than Theorem 1.11. Here, we treat the driving noises $(Z_{v})_{v\in V}$, which are generic noise processes, as part of the inputs to the equations (1.1). It turns out that the only technical assumptions that are essential are the $\delta$-robustness of the limiting configuration $(G,c,x,Z)$.
Proof of Theorem 1.11.
Step 1. We start by showing that the sequence $\{\mathcal{L}(G_{n},c^{n},x^{n},Z^{n},\underline{X}^{n},\underline{D}^{n})\}_{n\geq 1}$ is tight in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}^{3}])$. By the Skorokhod Representation Theorem, we can assume without loss of generality that $(G_{n},c^{n},x^{n},Z^{n})$ converges almost surely to $(G,c,x,Z)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$ as $n\to\infty$.
Fix any $\varepsilon\in(0,1)$. Since local convergence holds almost surely, we can find a compact subset $\mathcal{K}_{0}^{\varepsilon}$ of $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}]$ together with a sequence $(M_{m})_{m\geq 1}$ of positive real numbers such that
$\inf_{n}\mathbb{P}\Big[(G_{n},c^{n},x^{n},Z^{n})\in\mathcal{K}_{0}^{\varepsilon},\,\max_{v\in B_{G}(o,m)}\sum_{u\in N_{G}^{-}(v)}c_{uv}\leq M_{m}\Big]\geq 1-\frac{\varepsilon}{2}.$
@@
$\inf_{n}\mathbb{P}\Big[(G_{n},c^{n},x^{n},Z^{n})\in\mathcal{K}_{0}^{\varepsilon},\,\underline{L}_{v}^{n}\in\mathcal{M}(M_{m})\text{ and }\underline{D}_{v}^{n}\in\mathcal{M}(1),\,\forall v\in B_{G_{n}}(o_{n},m)\,\forall m\geq 1\Big]$
$\geq 1-\varepsilon,$
which implies the tightness of $\{\mathcal{L}(G_{n},c^{n},x^{n},Z^{n},\underline{L}^{n},\underline{D}^{n})\}_{n\geq 1}$ in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}^{3}])$ by Theorem A.2 and Lemma 2.4. As
$\underline{X}_{v}^{n}=x_{v}^{n}+Z_{v}^{n}-\underline{L}_{v}^{n},\quad v\in V_{n},$
we obtain that $\{\mathcal{L}(G_{n},c^{n},x^{n},Z^{n},\underline{X}^{n},\underline{D}^{n})\}_{n\geq 1}$ is tight in $\mathcal{P}(\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}^{3}])$ by combining Lemma 2.3 Lemma B.1 and [34, Lemma 12.7.3].
Step 2. By the Skorokhod Representation Theorem, we can find a $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}^{3}]$-valued random element $(G,c,x,Z,X,D)$ such that $(G_{n},c^{n},x^{n},Z^{n},\underline{X}^{n},\underline{D}^{n})\stackrel{{\scriptstyle a.s.}}{{\to}}(G,c,x,Z,X,D)$ in $\mathcal{G}_{*}[\mathbb{R}\times\mathcal{D}^{3}]$ as $n\to\infty$. In this step, we verify that $(X,D)$ is a solution to equations (1.1). As
$\underline{X}_{v}^{n}(t)$ $=x_{v}^{n}+Z_{v}^{n}(t)-\sum_{u\in N_{G_{n}}^{-}(v)}c_{uv}^{n}\mathbf{1}_{\{u\in D_{t}^{n}\}},\quad v\in V_{n},$
by Theorem B.2, for all $t\in[-1,\infty)$ such that $t$ is a continuity point of $(X_{v}(\cdot),\mathbf{1}_{\{v\in D_{\cdot}\}})$ for all $v\in V$ (the set of such $t$ is co-countable and thus dense in $1,\infty)$), it holds that
$X_{v}(t)$ $=x_{v}+Z_{v}(t)-\sum_{u\in N_{G}^{-}(v)}c_{uv}\mathbf{1}_{\{u\in D_{t}\}},\quad v\in V.$
@@
by Corollary B.3. To prove $\tau_{v}=\lim_{n\to\infty}\underline{\tau}_{v}^{n}$, we can use the same argument as in the proof of [15, Proposition 5.8], as the limit process $X_{v}$ satisfies the downward crossing property.